# Uncertainty Quantification with the Empirical Neural Tangent Kernel

### v1

@@ 5.3 Image Classification - Uncertainty
In Figure 2, we compare the UQ performance of NUQLS, DE, SWAG, LLA* (LLA with a last-layer and KFAC approximation), and MC-Dropout, on larger image classification tasks. While previous works have measured the uncertainty quantification performance through entropy, ECE, and AUC-ROC, we propose a new metric. The previous measurements are all based on calibration of the softmax output for classification. However, the softmax probabilities are not a notion of uncertainty, but are instead a point prediction of class probability; uncertainty should accompany a point prediction, and quantify the spread of possible values. Instead, we propose to use the variance of the maximum predicted softmax probability (VMSP), for a given test point, as a better quantifier of uncertainty.
Figure 2 presents a violin plot of the VMSP for three test-groups: correctly predicted in-distribution (FashionMNIST, CIFAR-10) test points, incorrectly predicted in-distribution test points, and out-of distribution (MNIST, CIFAR-100) test points. We would expect that there should be smaller uncertainty for ID test points that have been correctly predicted, and larger uncertainty for incorrectly predicted ID and OoD test points. We compare against a completely randomized baseline method (BASE), where we sample $10$ standard normal realizations of logits, passed through a softmax. In Table 7 in Section D.6, we display the corresponding median values for each method in each test group. Ideally, a method should far outperform the baseline, so we can use the median difference between a method and the baseline as a way to compare different methods. We see that NUQLS performs as well as or better than all other methods, including the SOTA method DE. Furthermore, we would like the ID correct distribution to have a positive-skew, while ID incorrect and OoD distributions should have a negative-skew. We also display the sample skew values in Table 7. In this metric, our method again performs as well as or better than all other methods.
5.4 Image Classification - Predictive
We now evaluate the peformance of the novel predictive mean on image classification tasks. We compare NUQLS against the MAP NN, DE, MC-Dropout, SWAG, LLA* and VaLLA, on the MNIST and FashionMNIST datasets, using the LeNet5 network. We compare test cross-entropy (NLL), test accuracy (ACC), ECE, and the AUC-ROC measurement for maximum softmax probability (AUC-ROC) and entropy (OOD-AUC) as the detector. The last two metrics evaluate a methods ability to detect out-of distribution points. We display the results in Table 2. We see that NUQLS performs the best in ECE, AUC-ROC and OOD-AUC, and is competitive in the other metrics, while having the second fastest wall-time.
@@ 6 Conclusion
We have presented NUQLS, a Bayesian, post-hoc UQ method that approximates the predictive distribution of an over-parameterized NN through GD/SGD training of linear networks, allowing scalability without sacrificing performance. Under assumptions on the loss function, this predictive distribution reduces to a GP using the NTK. We find that our method is competitive with, and often far outperforms, existing UQ methods on regression and classification tasks, whilst providing a novel connection between NNs, GPs and the NTK. Potential future research could involve extending Theorem 3.2 to broader classes of loss functions and other deterministic and stochastic optimization algorithms.
Acknowledgements
Fred Roosta was partially supported by the Australian Research Council through an Industrial Transformation Training Centre for Information Resilience (IC200100022).
@@ Appendix A Proofs
Proof of Lemma 3.1.
First, we note that by the assumption on $\ell$, a solution to 5 always exists. Suppose to the contrary that 5 has two distinct solutions $\tilde{\bm{\theta}}$ and $\widehat{\bm{\theta}}$ such that $\tilde{\bm{\theta}}\neq\widehat{\bm{\theta}}$. Since $\tilde{\bm{\theta}}\in{\textnormal{Range}}([{\mathbf{J}}(\widehat{\bm{\theta}},\mathcal{X})]^{{{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}}})$ and $\widehat{\bm{\theta}}\in{\textnormal{Range}}([{\mathbf{J}}(\widehat{\bm{\theta}},\mathcal{X})]^{{{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}}})$, i.e., $(\tilde{\bm{\theta}}-\widehat{\bm{\theta}})\perp{\textnormal{Null}}([{\mathbf{J}}(\widehat{\bm{\theta}},\mathcal{X})])$, it follows that ${\mathbf{J}}(\widehat{\bm{\theta}},\mathcal{X})\tilde{\bm{\theta}}\neq{\mathbf{J}}(\widehat{\bm{\theta}},\mathcal{X})\widehat{\bm{\theta}}$, which in particular implies $\langle\nabla{\bf f}(\widehat{\bm{\theta}},{\bf x}_{i}),\tilde{\bm{\theta}}\rangle\neq\langle\nabla{\bf f}(\widehat{\bm{\theta}},{\bf x}_{i}),\widehat{\bm{\theta}}\rangle$ for all $i=1,\ldots,n$. Consider $\bar{\bm{\theta}}=(\tilde{\bm{\theta}}+\widehat{\bm{\theta}})/2$. By strict convexity on $\ell$, we have
@@
    ${\mathbf{J}}(\widehat{\bm{\theta}},\mathcal{X})\triangleq\begin{bmatrix}{\mathbf{J}}^{{{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}}}(\widehat{\bm{\theta}},{\bf x}_{1})\;\ldots\;{\mathbf{J}}^{{{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}}}(\widehat{\bm{\theta}},{\bf x}_{n})\end{bmatrix}^{{{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}}}\in^{nc\times p},$
@@
    where ${\bf v}\in{\textnormal{Range}}\left(\left[{\mathbf{J}}(\widehat{\bm{\theta}},\mathcal{X})\right]^{{{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}}}\right)$. In particular, for ${\bf z}={\bf 0}$, by Lemma 3.1, we must have that ${\bf v}=\bm{\theta}^{\ddagger}$, where $\bm{\theta}^{\ddagger}$ is the solution to 5. Therefore,
@@
    Let $\bm{\theta}^{\star}$ be any solution to 3. The full row-rank assumption on ${\mathbf{J}}(\widehat{\bm{\theta}},\mathcal{X})$ implies that we must have ${\bf g}(\bm{\theta}^{\star},\mathcal{X})={\bf 0}$, i.e., $\nabla\ell(\widetilde{{\bf f}}(\bm{\theta}^{\star},{\bf x}_{i}),{\bf y}_{i})={\bf 0}$, $i=1,\ldots,n$. By the $\mu$-strong convexity assumption on (.$\ell(.,{\bf y})$ with respect to its first argument, it is easy to see that $\mathcal{L}(\bm{\theta})$ satisfies the Polyak-Łojasiewicz inequality (Karimi et al., 2016) with constant $2\mu\lambda$ where
@@
    and $\sigma_{\min}({\mathbf{J}}(\widehat{\bm{\theta}},{\bf x}_{i}))$ is the smallest non-zero singular value of ${\mathbf{J}}(\widehat{\bm{\theta}},{\bf x}_{i})$. Indeed, let $\bm{\theta}^{\star}$ be any solution to 3. From $\mu$-strong convexity of (.$\ell(.,{\bf y})$ with respect to its first argument, for any $\bm{\theta}$, we have
@@
    for some contact $10<\rho<$ 1. This in particular implies that for any $\epsilon>0$,
@@ Appendix B Related Works: Further Details and Discussions
NUQLS shares notable similarities with, and exhibits distinct differences from, several prior works, which are briefly discussed below.
@@
NUQLS can be seen as an extension of the “sample-then-optimize” framework for posterior sampling of large, finite-width NNs (Matthews et al., 2017). In this context, the work of Antorán et al. (2022), henceforth referred to as Sampling-LLA, enables drawing samples from the posterior distribution of the LLA in a manner analogous to Algorithm 1. In this approach, a series of regularized least-squares regression problems are constructed, and the collection of their solutions is shown to be distributed according to the LLA posterior. An EM algorithm is then employed for hyperparameter tuning. In addition to the fundamental differences between NUQLS and LLA-inspired methods mentioned earlier, Sampling-LLA has notable distinctions from Algorithm 1. First, the objective functions in Sampling-LLA have non-trivial minimum values. As a result, the convergence of SGD for such problems necessitates either a diminishing learning rate (Bubeck et al., 2015), which slows down convergence, or the adoption of variance reduction techniques (Roux et al., 2012; Shalev-Shwartz & Zhang, 2013; Johnson & Zhang, 2013), which can introduce additional computational and memory overhead. By contrast, in overparameterized settings and under the assumptions of Theorem 3.2, the optimization problem in 3 allows interpolation. Consequently, SGD can employ a constant step size for convergence, improving optimization efficiency (Garrigos & Gower, 2023). Second, the inherent properties of the LLA framework, which require a positive definite Hessian or its approximation, necessitate regularizing the least-squares term in the subproblem of Sampling-LLA. This results in a strongly convex problem with a unique solution. Consequently, to generate a collection of solutions, Sampling-LLA constructs a random set of such subproblems, each involving fitting the linearized network to random outputs. These random outputs are sampled from a zero-mean Gaussian, with covariance given by the Hessian of the loss function, evaluated on the data. In contrast, the subproblem of NUQLS, i.e., 3, involves directly fitting the training data, and the ensemble of solutions is constructed as a result of random initialization of the optimization algorithm. Hence, the uncertainty captured by NUQLS arises naturally from the variance of solutions in the overparameterized regime, without the need for additional regularization or artificially constructed subproblems.
@@
While the Sampling-LLA method enhances the scalability of LLA, the competing method Variational LLA (VaLLA) (Ortega et al., 2023) offers comparable or superior UQ performance while significantly reducing computation time. VaLLA achieves this by computing the LLA predictive distribution using a variational sparse GP with an NTK. Another competing LLA extension is Accelerated LLA (ELLA), which uses a Nyström approximation of the functional LLA covariance matrix (Deng et al., 2022), and seems to attain similar performance to VaLLA, again at a reduced cost compared to Sampling-LLA. We compare the performance of NUQLS to Sampling-LLA, VaLLA and ELLA in Section D.2.
B.2 Ensemble Framework.
@@
In Lee et al. (2019), infinitely wide neural networks were shown to follow a GP distribution; however, this GP did not correspond to a true predictive distribution. Building on this, He et al. (2020) introduced a random, untrainable function to an infinitely wide neural network, deriving a GP predictive distribution using the infinite-width NTK. An ensemble of these modified NNs was then interpreted as samples from the GP’s predictive posterior. In contrast, our method demonstrates that trained, finite-width, unmodified linearized networks are inherently samples from a GP with an NTK kernel. While their method, Bayesian Deep Ensembles (BDE), shares some conceptual similarities with ours, we omit it from our experiments for several reasons. Firstly, while it performed well in a toy regression problem in the original paper, it performed poorly on the alternate toy regression problem found in Figure 4 in Section D.3; this is likely due to the infinite-width NTK being a poor approximation in practice. Second, the method is computationally more expensive than DE, particularly for classification tasks, where it requires additional parameter tuning on a validation set. Since our goal is to provide a UQ method that is more computationally efficient than the state-of-the-art SOTA method DE while maintaining competitive performance, BDE is not considered a suitable candidate for comparison.
@@
In Madras et al. (2019), the authors propose a method called “local ensembles”, which perturbs the parameters of a trained network along directions of small loss curvature to create an ensemble of nearly loss-invariant networks. The uncertainty of the original network is then quantified as the standard deviation across predictions in the ensemble. Building on this approach, Miani et al. (2024), in a method called Sketched Lanczos Uncertainty (SLU), use the GGN approximation of the Hessian of the loss to identify these directions and introduce a sketched Lanczos algorithm (Meurant, 2006) to efficiently compute them. We compare the performance of NUQLS to SLU in Section D.1.

### v2

@@ 4.4 Image Classification - Uncertainty
Figure 3: Violin plot of VMSP, for correctly predicted ID test points, incorrectly predicted ID test points, and OoD test points. Median is shown, with violin width depicting density. Low variance is expected for ID correct points, and large variance for ID incorrect and OoD points.
We now compare the UQ performance of NUQLS, DE, SWAG, LLA* (LLA with a last-layer and KFAC approximation), and MC-Dropout (MC), on larger image classification tasks. We take variance of the maximum predicted softmax probability (VMSP), for a given test point, as the correct quantifier of uncertainty in this setting (see Appendix˜C for justification).
Figure˜3 presents a violin plot of the VMSP for three test-groups: correctly predicted in-distribution (FashionMNIST, CIFAR-10) test points, incorrectly predicted in-distribution test points, and out-of distribution (MNIST, CIFAR-100) test points. We would expect that there should be, on average, much smaller uncertainty for ID test points that have been correctly predicted, and larger uncertainty for incorrectly predicted ID and OoD test points. We compare against a completely randomized baseline method (BASE), where we sample $10$ standard normal realizations of logits, passed through a softmax. In Table˜10 in Section˜H.8, we display the corresponding median and sample skew values for each method in each test group, to quantify the distribution of VMSP for each test set. Ideally, a method should far outperform the baseline, so we can use the median and sample skew difference between a method and the baseline as a way to compare different methods. We see that NUQLS outperforms all other methods, including the SOTA method DE. In Section˜H.7 we provide additional experimental evaluation of NUQLS. In Figure˜6 we evaluate NUQLS on a ResNet50 trained on both SVHN and ImageNet, displaying the scalability of our method, as well as providing comparison with other competing methods, including Bayesian Deep Ensembles (BDE), Spectral-Normalized Neural Gaussian Process (SNGP), BatchEnsemble (BE), and Stochastic Gradient Langevin Dynamics (SGLD). Against these competing methods, NUQLS performs the strongest.
@@ 5 Conclusion
We have presented NUQLS, a Bayesian, post-hoc UQ method that approximates the predictive distribution of an over-parameterized NN through GD/SGD training of linear networks, allowing scalability without sacrificing performance. Under assumptions on the loss function, this predictive distribution reduces to a GP using the NTK. We find that our method is competitive with, and often far outperforms, existing UQ methods on regression and classification tasks, whilst providing a novel connection between NNs, GPs and the NTK.
Limitation.
A theoretical limitation of this work is that its connection to the NTK-GP does not extend to loss functions that violate the assumptions of Theorem˜3.2, such as the cross-entropy loss. The strong empirical performance of NUQLS on classification tasks motivates future research to extend Theorem˜3.2 to broader classes of loss functions and alternative optimization algorithms. A further limitation of the method is the dependence of NUQLS on the linearization approximation. As can be seen in Section˜H.1, when the target neural network is poorly trained, and hence the loss-landscape is far from flat around the ’trained’ parameters, the performance of NUQLS suffers. However, when the network is well-trained, as is common in practical settings, the linear approximation holds and hence NUQLS performs well.
Acknowledgments and Disclosure of Funding
Liam Hodgkinson is supported by the Australian Research Council through a Discovery Early Career Researcher Award (DE240100144). Fred Roosta was partially supported by the Australian Research Council through an Industrial Transformation Training Centre for Information Resilience (IC200100022).
@@ Appendix A Proofs
Proof of Lemma˜3.1.
First, we note that by the assumption on $\ell$, a solution to ˜5 always exists. Suppose to the contrary that ˜5 has two distinct solutions $\tilde{\bm{\theta}}$ and $\widehat{\bm{\theta}}$ such that $\tilde{\bm{\theta}}\neq\widehat{\bm{\theta}}$. Since $\tilde{\bm{\theta}}\in{\textnormal{Range}}([{\mathbf{J}}(\widehat{\bm{\theta}},\mathcal{X})]^{{{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}}})$ and $\widehat{\bm{\theta}}\in{\textnormal{Range}}([{\mathbf{J}}(\widehat{\bm{\theta}},\mathcal{X})]^{{{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}}})$, i.e., $(\tilde{\bm{\theta}}-\widehat{\bm{\theta}})\perp{\textnormal{Null}}([{\mathbf{J}}(\widehat{\bm{\theta}},\mathcal{X})])$, it follows that $^{\mathbf{J}}(\widehat{\bm{\theta}},\mathcal{X})\tilde{\bm{\theta}}\neq{\mathbf{J}}(\widehat{\bm{\theta}},\mathcal{X})\widehat{\bm{\theta}}$, which in particular implies $\langle\nabla{\bf f}(\widehat{\bm{\theta}},{\bf x}_{i}),\tilde{\bm{\theta}}\rangle\neq\langle\nabla{\bf f}(\widehat{\bm{\theta}},{\bf x}_{i}),\widehat{\bm{\theta}}\rangle$ for all $i=1,\ldots,n$. Consider $\bar{\bm{\theta}}=(\tilde{\bm{\theta}}+\widehat{\bm{\theta}})/2$. By strict convexity on $\ell$, we have
@@
[…]
@@
    where ${\bf v}\in{\textnormal{Range}}\left(\left[{\mathbf{J}}(\widehat{\bm{\theta}},\mathcal{X})\right]^{{{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}}}\right)$. In particular, for ${\bf z}={\bf 0}$, by Lemma˜3.1, we must have that ${\bf v}=\bm{\theta}^{\ddagger}$, where $\bm{\theta}^{\ddagger}$ is the solution to ˜5. Therefore,
@@
    Let $\bm{\theta}^{\star}$ be any solution to ˜3. The full row-rank assumption on ${\mathbf{J}}(\widehat{\bm{\theta}},\mathcal{X})$ implies that we must have ${\bf g}(\bm{\theta}^{\star},\mathcal{X})={\bf 0}$, i.e., $\nabla\ell(\widetilde{{\bf f}}(\bm{\theta}^{\star},{\bf x}_{i}),{\bf y}_{i})={\bf 0}$, $i=1,\ldots,n$. By the $\mu$-strong convexity assumption on $\ell(.,{\bf y})$ with respect to its first argument, it is easy to see that $\mathcal{L}(\bm{\theta})$ satisfies the Polyak-Łojasiewicz inequality (Karimi et al., 2016) with constant $2\mu\lambda$ where
@@
    and $\sigma_{\min}({\mathbf{J}}(\widehat{\bm{\theta}},{\bf x}_{i}))$ is the smallest non-zero singular value of ${\mathbf{J}}(\widehat{\bm{\theta}},{\bf x}_{i})$. Indeed, let $\bm{\theta}^{\star}$ be any solution to ˜3. From $\mu$-strong convexity of $\ell(.,{\bf y})$ with respect to its first argument, for any $\bm{\theta}$, we have
@@
    for some contact $0<\rho<1$. This in particular implies that for any $\epsilon>0$,
@@ Appendix B Related Works: Further Details and Discussions
NUQLS shares notable similarities with, and exhibits distinct differences from, several prior works, which are discussed in-depth below.
@@
NUQLS can be seen as an extension of the “sample-then-optimize” framework for posterior sampling of large, finite-width NNs (Matthews et al., 2017). In this context, the work of Antorán et al. (2022), henceforth referred to as Sampling-LLA, enables drawing samples from the posterior distribution of the LLA in a manner analogous to Algorithm˜1. In this approach, a series of regularized least-squares regression problems are constructed, and the collection of their solutions is shown to be distributed according to the LLA posterior. An EM algorithm is then employed for hyperparameter tuning. In addition to the fundamental differences between NUQLS and LLA-inspired methods mentioned earlier, Sampling-LLA has notable distinctions from Algorithm˜1. First, the objective functions in Sampling-LLA have non-trivial minimum values. As a result, the convergence of SGD for such problems necessitates either a diminishing learning rate (Bubeck et al., 2015), which slows down convergence, or the adoption of variance reduction techniques (Roux et al., 2012; Shalev-Shwartz & Zhang, 2013; Johnson & Zhang, 2013), which can introduce additional computational and memory overhead. By contrast, in overparameterized settings and under the assumptions of Theorem˜3.2, the optimization problem in ˜3 allows interpolation. Consequently, SGD can employ a constant step size for convergence, improving optimization efficiency (Garrigos & Gower, 2023). Second, the inherent properties of the LLA framework, which require a positive definite Hessian or its approximation, necessitate regularizing the least-squares term in the subproblem of Sampling-LLA. This results in a strongly convex problem with a unique solution. Consequently, to generate a collection of solutions, Sampling-LLA constructs a random set of such subproblems, each involving fitting the linearized network to random outputs. These random outputs are sampled from a zero-mean Gaussian, with covariance given by the Hessian of the loss function, evaluated on the data. In contrast, the subproblem of NUQLS, i.e., ˜3, involves directly fitting the training data, and the ensemble of solutions is constructed as a result of random initialization of the optimization algorithm. Hence, the uncertainty captured by NUQLS arises naturally from the variance of solutions in the overparameterized regime, without the need for additional regularization or artificially constructed subproblems.
@@
While the Sampling-LLA method enhances the scalability of LLA, the competing method Variational LLA (VaLLA) (Ortega et al., 2023) offers comparable or superior UQ performance while significantly reducing computation time. VaLLA achieves this by computing the LLA predictive distribution using a variational sparse GP with an NTK. Another competing LLA extension is Accelerated LLA (ELLA), which uses a Nyström approximation of the functional LLA covariance matrix (Deng et al., 2022), and seems to attain similar performance to VaLLA, again at a reduced cost compared to Sampling-LLA. We compare the performance of NUQLS to Sampling-LLA, VaLLA and ELLA in Section˜H.3.
Note that work by (Miani et al., 2024b) approximates an LLA by taking the covariance as the projection onto the null space of the GGN matrix, in order to compute a posterior that retains the same performance as the original network on training data. To compute samples, this work uses alternating projections.
B.2 SNGP
We briefly describe the Spectral-Normalized Neural Gaussian Process (SNGP). SNGP lies at the intersection of feature-space methods and Bayesian methods, and similarly to NUQLS combines a GP with a NN. Specifically, SNGP adds a weight normalization step during training, and then replaces the output layer of NN with a GP that takes the feature extractor of the network as an input. While this method is also not post-hoc, and is not strictly a Bayesian method, we still compare SNGP with NUQLS in Figure˜6 and Table˜11, where we observe superior performance of NUQLS.
B.3 Ensemble Framework.
@@ B.2 SNGP
In Lee et al. (2019), infinitely wide neural networks were shown to follow a GP distribution; however, this GP did not correspond to a true predictive distribution. Building on this, He et al. (2020) introduced a random, untrainable function to an infinitely wide neural network, deriving a GP predictive distribution using the infinite-width NTK. An ensemble of these modified NNs was then interpreted as samples from the GP’s predictive posterior. In contrast, our method demonstrates that trained, finite-width, unmodified linearized networks are inherently samples from a GP with an NTK kernel. While their method, Bayesian Deep Ensembles (BDE), shares some conceptual similarities with ours, we omit it from our main experiments for several reasons. Firstly, the posterior analysis of BDE is valid only in the limit of large model sizes. For smaller datasets and models, the infinite-width NTK differs greatly from the empirical NTK (see Fort et al. (2020)). We can see that in this regime, in both Figure˜4 and Table˜8, that NUQLS outperforms BDE. Secondly, the method is computationally more expensive than DE, due to the need to compute the untrainable function ($\delta(.)$) and tune the scaling hyperparameter for classification. This contradicts our goal to provide a UQ method that is computationally more efficient than the state-of-the-art (SOTA) method DE while maintaining competitive performance. Finally, for larger model sizes, we provide comparison of NUQLS to BDE in Figure˜6 and Table˜11, where we observe that NUQLS outperforms BDE.
@@
In Madras et al. (2019), the authors propose a method called “local ensembles”, which perturbs the parameters of a trained network along directions of small loss curvature to create an ensemble of nearly loss-invariant networks. The uncertainty of the original network is then quantified as the standard deviation across predictions in the ensemble. Building on this approach, Miani et al. (2024a), in a method called Sketched Lanczos Uncertainty (SLU), use the GGN approximation of the Hessian of the loss to identify these directions and introduce a sketched Lanczos algorithm (Meurant, 2006) to efficiently compute them. We compare the performance of NUQLS to SLU in Section˜H.2.